OpenStudy (anonymous):

A battery can operate 8 toys for 21 hours. For how long could the battery operate 15 bears?

OpenStudy (anonymous):

I was hope you would answer

OpenStudy (anonymous):

I have been messing arounf with it and I do have an answer

OpenStudy (anonymous):

i done 8*21=15x where x is the hours

OpenStudy (callisto):

@Hollywood_chrissy I was doing in the same way as you :S

OpenStudy (anonymous):

:) but I dont stand what I done.

OpenStudy (diyadiya):

A battery can operate 8 toys for 21 hours so for 1 toy =21/8 hours 15 toys =21/8 * 15 ?

OpenStudy (anonymous):

Less toys more time so it's inversely proportional. Using unitary method: 8 toys for 21 hours 1 toys for 21*8 hours 15 toys in \( (21\times8)/15 = \frac{56}{5} = 11.2 \) hours

OpenStudy (diyadiya):

Isn't 1 toy 21/8 hrs?

OpenStudy (anonymous):

If we use batteries of 8 toys in just 1 toy then the battery will last long. So more time.

OpenStudy (diyadiya):

OOOOh okay!!

OpenStudy (anonymous):

NO worries diya, I got deceived too :P

OpenStudy (callisto):

Here comes the question...Is it 8 toys for 21 hours each (total time = 8x12 hours )or 21 hours for 8 toys ... That's what I'm thinking of.... :S (personally i think it is the first case)

OpenStudy (callisto):

This question is somehow like the previous one you asked..

OpenStudy (anonymous):

@Hollywood_chrissy: Do you know the right answer? :)

OpenStudy (kropot72):

Hollywood_chrissy and FoolForMath both have the correct answer. Increasing the number of toys will drain the single battery faster.

OpenStudy (anonymous):

can I get an good answer????

OpenStudy (callisto):

Assume you've got the correct numerical answer. Here is the explanation. The total operating hours for the battery = 8x21 = 168 hours, understand this part?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Now, you have the same battery, that means the total operating hours for the battery remains unchanged, that is 168 hours. understand this part?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Then, you have 15 bears now. You want to use the battery to operate these 15 toys (equally), understand this part?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So let x be the number of operating hours per bear. number of bear * number of operating hours per bear = total number of working hours 15 x = 168 Understand this part?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

solve x and you'll get the answer you want. By the way, is the answer correct?

OpenStudy (anonymous):

yes 11.2 hours

OpenStudy (callisto):

Good :) Do you understand it now?

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

thank you once more.

OpenStudy (callisto):

Glad to hear :) and you're welcome :)